The density of an object is W/V. If the density of an object is 5.25x10^-7 oz/ft^3?
and the volume is 2.2x10^3 ft^3, then find the weight?
Honestly i get to an answer 2.39 x10^-10 but i guess it's nor right
So i'm be happy if somebody can help me out?

Answer by jim_thompson5910(35256)

D = W/V

5.25x10^-7 = W/(2.2x10^3)

(5.25x10^-7)*(2.2x10^3) = W

(5.25*2.2)x(10^-7*10^3) = W

(11.55)x(10^-4) = W

(1.155 x 10^1)x(10^-4) = W

1.155 x (10^1*10^-4) = W

1.155 x 10^-3 = W

W = 1.155 x 10^-3

So the weight is roughly 1.155 x 10^-3 ounces or 0.001155 ounces.